A building contractor estimated labor cost for a job will be $25,000 more than the material cost. Tha total labor and material cost is $95,000
a. What is the estimated material cost?
b. What is the estimated labor cost?
I have (m = material) 2m + $25,000 = $95,000.
To me this set up does not look correct to me.

Answer by richwmiller(17219)

Why 2m?
l+m=95000
l=m+25000
substitute
m+m+25000=95000
2m+25000=95000
What is wrong with that?
2m=70k
m=35k
l=35k+25k
l=60k